### v1

@@ 3 Concluding Remarks
Stable Attractiveness $1.081\times 10^{-7}$ 0.063 [100, 1] [512, 256]
Diffusion Dominance $1.021\times 10^{-7}$ 0.270 [200, 1] [1024, 512]
ver1.5 Trustworthiness $1.166\times 10^{-7}$ 0.193 [50, 1] [512, 256]
Stable Attractiveness $1.018\times 10^{-7}$ 0.083 [100, 1] [1024, 512]
Diffusion Dominance $1.019\times 10^{-7}$ 0.169 [50, 1] [512, 256]
ver 2.1 Trustworthiness $1.012\times 10^{-7}$ 0.052 [200, 1] [256, 128]
@@
To estimate the average potential outcome for each facial trait, we applied the methodology described earlier, using 2-fold cross-fitting in combination with the selected hyperparameters to estimate both the deconfounder and outcome models. All optimization settings — including the choice of optimizer, batch size, number of training epochs, gradient clipping, and early stopping — were held consistent with the prior application for comparability. We then compared our results to the linear regression results with a discretized treatment and the same covariates, which mirrored the original paper’s analysis in Lindholm et al., (2024).
@@
$Y_{jj^{\prime}(i)}=\begin{cases}0&\text{if respondent $i$ answers that argument $j$ is more persuasive}\\ 1&\text{if respondent $i$ answers that both arguments are equally persuasive}\\ 2&\text{if respondent $i$ answers that argument $j^{\prime}$ is more persuasive}\end{cases}$
@@
    Use the entire data to solve the optimization problem given in Equation (S11) and simultaneously obtain $\hat{\bm{f}}(\bm{R}_{j}):=\bm{f}(\bm{R}_{j};\hat{\bm{\lambda}})$ and $\hat{\mu}(T_{j},\bm{R}_{j}):=\mu(T_{j},\bm{f}(\bm{R}_{j};\hat{\bm{\lambda}});\hat{\bm{\theta}})$
@@
Appendix S4 Proof of Proposition 1

### v2

@@ 3 Concluding Remarks
Stable Attractiveness $1.027\times 10^{-7}$ 0.147 [50, 1] [256, 128]
Diffusion Dominance $1.364\times 10^{-7}$ 0.079 [50, 1] [512, 256]
ver1.5 Trustworthiness $1.001\times 10^{-7}$ 0.124 [100, 1] [256, 128]
Stable Attractiveness $1.006\times 10^{-7}$ 0.234 [50, 1] [256, 128]
Diffusion Dominance $1.018\times 10^{-7}$ 0.069 [50, 1] [512, 256]
ver 2.1 Trustworthiness $1.266\times 10^{-7}$ 0.245 [50, 1] [512, 256]
@@
To estimate the average potential outcome for each facial trait, we applied the methodology described earlier, using 2-fold cross-fitting in combination with the selected hyperparameters to estimate both the deconfounder and outcome models. All optimization settings — including the choice of optimizer, batch size, number of training epochs, gradient clipping, and early stopping — were held consistent with the prior application for comparability. We also combined the estimates from the two different generative models (Stable Diffusion versions 1.5 and 2.1) by the optimal linear combination method proposed in the previous section. We then compared our results to the linear regression results with a discretized treatment and the same covariates, which mirrored the original paper’s analysis in Lindholm et al., (2024). All the results are presented in Figure S4.
Figure S4: Estimated predictive effects of three facial traits (Attractiveness, Dominance, and Trustworthiness) using GPI with Stable Diffusion (SD) version 1.5 (top) and version 2.1 (second), their optimal linear combination (third), and OLS (bottom). In each panel, solid and dash lines represent estimates with and without covariate adjustment, respectively. Blue and red shaded areas indicate the corresponding 95% confidence intervals.
@@
[…]
@@
[…]
@@
Appendix S4 Proofs
S4.1 Proof of Proposition 1
@@
S4.2 Proof of Proposition 2
Proof.
By Theorem 1, we know that each estimator $\hat{\tau}_{j}$ ($j=1,2$) has the influence function representation:
$\sqrt{N}(\hat{\tau}_{j}-\tau)=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\psi_{j}(\mathcal{D}_{ij})+o_{P}(1)$
where $\psi_{j}(\mathcal{D}_{ij})$ is the influence function for the estimator $\hat{\tau}_{j}$ at the observation $\mathcal{D}_{ij}$. Therefore, for any weight $\omega$, the estimator $\hat{\tau}=\omega\hat{\tau}_{1}+(1-\omega)\hat{\tau}_{2}$ has the influence function representation:
$\sqrt{N}(\hat{\tau}-\tau)=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\{\omega\psi_{1}(\mathcal{D}_{i1})+(1-\omega)\psi_{2}(\mathcal{D}_{i2})\}+o_{P}(1)$
Therefore, by central limit theorem, for any $\omega$, we have
$\sqrt{N}(\hat{\tau}-\tau)\xrightarrow{d}\mathcal{N}(0,V(\omega))$
where $V(\omega)=\mathbb{E}[(\omega\psi_{1}(\mathcal{D}_{i1})+(1-\omega)\psi_{2}(\mathcal{D}_{i2}))^{2}]$.
Then, since $V(\omega)$ is a quadratic function of $\omega$, we can minimize the asymptotic variance $V(\omega)$ with respect to $\omega$ as follows,
$V(\omega)$ $=\omega^{2}\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})^{2}]+(1-\omega)^{2}\mathbb{E}[\psi_{2}(\mathcal{D}_{i2})^{2}]+2\omega(1-\omega)\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})\psi_{2}(\mathcal{D}_{i2})]$
$=\{\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})^{2}]+\mathbb{E}[\psi_{2}(\mathcal{D}_{i2})^{2}]-2\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})\psi_{2}(\mathcal{D}_{i2})]\}\omega^{2}$
$\quad\quad+2\{\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})\psi_{2}(\mathcal{D}_{i2})]-\mathbb{E}[\psi_{2}(\mathcal{D}_{i2})^{2}]\}\omega+\mathbb{E}[\psi_{2}(\mathcal{D}_{i2})^{2}].$
As it is a quadratic function of $\omega$, it is minimized at
$\omega^{*}=\frac{\mathbb{E}[\psi_{2}(\mathcal{D}_{i2})^{2}]-\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})\psi_{2}(\mathcal{D}_{i2})]}{\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})^{2}]+\mathbb{E}[\psi_{2}(\mathcal{D}_{i2})^{2}]-2\mathbb{E}[\psi_{1}(\mathcal{D}_{i1})\psi_{2}(\mathcal{D}_{i2})]}$